OpenStudy (anonymous):

show all work. Add the following polynomials: (1/4x^2-1/6x- 3/8)+(1/4x^2-1/4x-1/2)

OpenStudy (anonymous):

which i have 1/4x^2 - 4/24x - 3/8 1/4x^2 - 6/24x - 4/8 am i doing this right

OpenStudy (anonymous):

1/4x^2 + 1/4x^2 = 1/2 x^2 -1/6x + - 1/4x = -2/12x - 3/12x = -5/12x -3/8 + - 1/2 = -3/8 - 4/8 = -7/8x comes down to: 1/2x^2 - 5/12x - 7/8x